OpenStudy (anonymous):

How to show that x=3 is the only real solution of the equation x^3-4x-15=0?

OpenStudy (anonymous):

x^3 -3x^2 +3x^2 -9x + 5x -15=0 x^2(x-3) + 3x(x-3) +5(x-3)=0 (x-3) (x^2+3x+5)=0

OpenStudy (anonymous):

Now, show that second quadratic equation has no real solution HINT: show that b^2-4ac<0